# Thomassen’s proof and Filippov’s proof of the Weak Jordan Theorem

### 19 Apr 2026 v3

@@ Abstract
Let us begin with a few necessary definitions. We denote by $I$ any compact real interval $I=[a,b]$ with $a<b$. If $f\colon I\to X$ is a map, we set $f[I]=\{f(x)\colon\;x\in I\}$, $f^{0}=\{f(x)\colon\;a<x<b\}$ and $\mathrm{ep}(f)=\{f(a),f(b)\}$. We also say that $f$ joins $f(a)$ to $f(b)$ in $X$. If $f(a)=f(b)$, we call $f$ closed. If $g\colon A\to B$ is any map and $C$ is any set, we define
@@
In 1992, C. Thomassen found in [17] an interesting proof of the JT based on planar graphs. He actually obtained combinatorial proofs of the more general Jordan–Schönflies theorem and of the theorem on classification of surfaces. In Section 2 we present only a part of [17], Thomassen’s proof of the WJT. It employs the following virtual configurations. A $K_{3,3}$-configuration has two disjoint three-element sets $U,V\subset\mathbb{R}^{2}$ and nine arcs $f_{u,v}\colon I\to\mathbb{R}^{2}$ for $u\in U$ and $v\in V$ such that $f_{u,v}$ joins $u$ to $v$ and
@@
for every two pairs $\langle u,v\rangle\neq\langle u^{\prime},v^{\prime}\rangle$. It is a plane drawing, without crossings, of the complete bipartite graph $K_{3,3}$; these configurations do not exist. A PL map is a map $f\colon[a,b]\to\mathbb{R}^{2}$ such that for a partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of $[a,b]$, all restrictions $f\,|\,[a_{i-1},a_{i}]$ for $i=1,2,\dots,k$ are non-constant linear maps. Every PL map is continuous. The straight segments $f[\,[a_{i-1},a_{i}]\,]$ ($\subset\mathbb{R}^{2}$) are the segments of $f[I]$. The points $f(a_{i})$ ($\in\mathbb{R}^{2}$) are the corners of $f[I]$. In a PL $K_{3,3}$-configuration, the nine arcs are PL maps. Thomassen’s proof of the WJT in [17], presented in Section 2, is the sequence claim 1 $\Rightarrow$ claim 2 $\Rightarrow$ claim 3 $\Rightarrow$ claim 4, where the four claims are as follows.
@@
    The JT holds for PL circuits.
@@
Another proof of the WJT is contained in the proof of the JT given by A. F. Filippov in [5] in 1950. We present this proof in Section 3. It can be divided in three steps.
@@
    such that $N(c)$ is the parity of the number of simple intersections of the vertical half-line going up from the point $c$ with $f[I]$ (we precisely define $N$ later). Then $N$ is constant.
@@
The first two auxiliary results concern closed (and non-compact) sets in metric spaces used in Filippov’s proof, such as vertical half-lines in $\mathbb{R}^{2}$. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is closed if the complement $M\setminus X$ is an open set, that is, if for every point $a\in M\setminus X$ there exists an $r>0$ such that the ball $B(a,r)\subset X\setminus M$.
@@
The second group of auxiliary results concerns compactness in metric spaces. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is compact if every sequence in $X$ has a convergent subsequence with a limit in $X$. We take for granted the compactness of $I=[a,b]$ ($\subset\mathbb{R}^{2}$).
@@
Proof. Suppose, for the contrary, that $f$ is not uniformly continuous. Then there exist $\varepsilon>0$ and sequences $(a_{n}),(b_{n})\subset M$ such that $\lim d(a_{n},b_{n})=0$, but $e(f(a_{n}),f(b_{n}))\geq\varepsilon$ for every $n$. Due to the compactness of $M$ we may assume that $\lim a_{n}=\lim b_{n}=a$ ($\in M$). But $\lim f(a_{n})=\lim f(b_{n})=f(a)$ ($\in N$) does not hold, in contradiction with the continuity of $f$ at $a$ $\Box$
@@
Let $X\subset\mathbb{R}^{2}$ be a compact set. Then there exist four points $u,v,r,s\in X$ such that $u_{x}$ and $v_{x}$ have, respectively, the maximum and minimum $x$-coordinate on $X$, and $r$ and $s$ have an analogous property in the $y$-coordinate.
@@ 2 Thomassen’s proof of the WJT
Claim 1. The JT holds for PL circuits. Suppose that $f\colon I\to\mathbb{R}^{2}$ is a PL circuit. We first prove that the open set $D=\mathbb{R}^{2}\setminus f[I]$ (Proposition 1.8) is disconnected: we define a continuous map $g\colon D\to\mathbb{R}$ such that $g[D]=\{0,1\}$. By Proposition 1.13, the set $D$ is disconnected because $\{0,1\}$ is a disconnected subset of $\mathbb{R}$.
@@
So it remains to prove that $L(f)$ and $R(f)$ are connected sets. The remaining properties stated in the JT, namely that one of the sets is bounded, the other is unbounded, and that $f[I]$ is the common boundary, follow easily.
We prove that $L(f)$ is connected (the argument for $R(f)$ is similar). We fix a point $c\in L(f)$ and for every point $d\in L(f)$ we define a connected set $C_{d}$ such that $c,d\in C_{d}\subset L(f)$. Then
@@
We fix one of the two orientations of $C[I]$ and consider the partition $D=\mathbb{R}^{2}\setminus C[I]=L(C)\cup R(C)$ of $D$ defined in Step 1. Let
@@
Each of these nine open sets is connected by Propositions 1.15 and 1.16 (the intersection graph spans a path), and they are mutually disjoint. Using Proposition 1.18, we obtain nine PL arcs $h_{i,j}\colon[a,b]\to U_{i,j}$ such that $h_{i,j}$ joins $\alpha_{i,j}=h_{i,j}(a)=g_{i,j}(a)$ ($\in\partial B_{i}$) to $\beta_{i,j}=h_{i,j}(b)=g_{i,j}(b)$ ($\in\partial B_{j}$) in $U_{i,j}$. We consider the first intersections of the straight segments $u_{i}\,\alpha_{i,j}$ and $u_{j}\,\beta_{i,j}$ with $h_{i,j}[I]$ (Proposition 1.9) and obtain nine PL arcs $p_{i,j}$ such that $p_{i,j}$ joins $u_{i}$ to $u_{j}$ and
@@ 3 Filippov’s proof of the WJT
We prove that the parity map $N$ is continuous on $S(f)$. Let $c\in S(f)$. We take a $\delta>0$ small enough such that the ball $B=B(c,\delta)\subset S(f)$ (Proposition 1.8) and that no vertical line that intersects $B$ contains a corner of $f[I]$ distinct from those whose preimages appear in $P(c)$. Let $c^{\prime}\in B$ with $c^{\prime}\neq c$. As we move $c$ to $c^{\prime}$, every simple component of $P(c)$ is preserved and every double component is either fixed or disappears or decays in two simple (point) components. It follows that $N(c)=N(c^{\prime})$. So $N$ is continuous at $c$ and on $S(f)$.
We obtain the main result of Step 1:
@@
Here is an outline of the proof of this equality: if $S_{i}$ is the set of simple components in $P(c,f_{i})$, $i=1,2,3$, then there is a bijection between
$S_{3}\,\text{ and }\,(S_{1}\times\{0\})\cup(S_{2}\times\{1\})\,.$
We leave details of the bijection to the reader as an exercise.
@@
Step 3. Some two points in the complement of any circuit are separated. Let $f\colon I\to\mathbb{R}^{2}$ be a circuit and $D=\mathbb{R}^{2}\setminus f[I]$. We may assume that $f[I]$ lies to the right of the $y$-axis. Let $l\in f[I]$ and $p\in f[I]$ be, respectively, a leftmost and a rightmost point in $f[I]$ (Propositions 1.6 1.19). These points divide $f$ in two arcs $f_{1},f_{2}\colon I\to\mathbb{R}^{2}$ such that $f_{1}^{0}\cap f_{2}^{0}=\emptyset$, $f_{1}[I]\cup f_{2}[I]=f[I]$ and $\mathrm{ep}(f_{1})=\mathrm{ep}(f_{2})=\{l,p\}$ (Proposition 1.21). Clearly, $l_{x}<p_{x}$ (Proposition 1.20). Let $\ell$ be any vertical line with $x$-coordinate in $(l_{x},p_{x})$. Let $a\in f[I]\cap\ell$ be the highest point of this intersection (Proposition 1.19). We may assume that $a\in f_{1}^{0}$. Similarly, let $b\in f_{1}^{0}\cap\ell$ be the lowest point of this intersection. We take a point $c\in\ell$ below $b$ but such that

### 22 Apr 2026 v4

@@ Abstract
JT, WJT and AT. Let us begin with a few necessary definitions. We denote by $I$ any compact real interval $I=[a,b]$ with $a<b$. If $f\colon I\to X$ is a map, we set $f[I]=\{f(x)\colon\;x\in I\}$, $f^{0}=\{f(x)\colon\;a<x<b\}$ and $\mathrm{ep}(f)=\{f(a),f(b)\}$. We also say that $f$ joins $f(a)$ to $f(b)$ in $X$. If $f(a)=f(b)$, we call $f$ closed. If $g\colon A\to B$ is any map and $C$ is any set, we define
@@
Thomassen’s proof of the WJT. In 1992, C. Thomassen found in [17] an interesting proof of the JT based on planar graphs. He actually obtained combinatorial proofs of the more general Jordan–Schönflies theorem and of the theorem on classification of surfaces. In Section 2 we present only a part of [17], Thomassen’s proof of the WJT. It employs the following virtual configurations. A $K_{3,3}$-configuration has two disjoint three-element sets $U,V\subset\mathbb{R}^{2}$ and nine arcs $f_{u,v}\colon I\to\mathbb{R}^{2}$ for $u\in U$ and $v\in V$ such that $f_{u,v}$ joins $u$ to $v$ and
@@
for every two pairs $\langle u,v\rangle\neq\langle u^{\prime},v^{\prime}\rangle$. It is a plane drawing, without crossings, of the complete bipartite graph $K_{3,3}$; these configurations do not exist. A PL map is a map $f\colon[a,b]\to\mathbb{R}^{2}$ such that for a partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of $[a,b]$, all restrictions $f\,|\,[a_{i-1},a_{i}]$ for $i=1,2,\dots,k$ are non-constant linear maps. Every PL map is continuous. The straight segments $f[\,[a_{i-1},a_{i}]\,]$ ($\subset\mathbb{R}^{2}$) are the segments of $f[I]$. The points $f(a_{i})$ ($\in\mathbb{R}^{2}$) are the corners of $f[I]$. In a PL $K_{3,3}$-configuration, the nine arcs are PL maps. Thomassen’s proof of the WJT in [17], that we present in Section 2, is the sequence claim 1 $\Rightarrow$ claim 2 $\Rightarrow$ claim 3 $\Rightarrow$ claim 4, where the four claims are as follows.
@@
    The Intermediate JT holds for PL circuits: the complement of any PL circuit is a disjoint union of two nonempty open connected sets.
@@
Filippov’s proof of the WJT. Another proof of the WJT is contained in the proof of the JT given by A. F. Filippov in [5] in 1950. We present this proof in Section 3. We divide it in three steps.
@@
    such that $N(c)$ is the parity of the number of simple intersections of the vertical half-line going up from the point $c$ with $f[I]$ (we provide a precise definition later). Then $N$ is constant.
@@
The first two auxiliary results concern closed sets in metric spaces. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is closed if the complement $M\setminus X$ is an open set, that is, if for every point $a\in M\setminus X$ there exists an $r>0$ such that the ball $B(a,r)\subset X\setminus M$. It is well known that $X\subset M$ is closed if and only if for every sequence $(a_{n})\subset X$ with $\lim a_{n}=a\in M$ we have $a\in X$.
@@
The use of the positivity of the distance between two disjoint plane sets, one of which is compact and the other is closed but non-compact (for example, a half-line), is an interesting maneuver in Filippov’s proof.
The second group of auxiliary results concerns compactness in metric spaces. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is compact if every sequence in $X$ has a convergent subsequence with a limit in $X$. We take for granted the compactness of intervals $I=[a,b]$ ($\subset\mathbb{R}^{2}$).
@@
Proof. Suppose, for the contrary, that $f$ is not uniformly continuous. Then there exist $\varepsilon>0$ and sequences $(a_{n}),(b_{n})\subset M$ such that $\lim d(a_{n},b_{n})=0$, but $e(f(a_{n}),f(b_{n}))\geq\varepsilon$ for every $n$. Due to the compactness of $M$, we may assume that $\lim a_{n}=\lim b_{n}=a$ ($\in M$). But $\lim f(a_{n})=\lim f(b_{n})=f(a)$ ($\in N$) does not hold, in contradiction with the continuity of $f$ at $a$. $\Box$
@@
Let $X\subset\mathbb{R}^{2}$ be a compact set. Then there exist four points $u,v,r,s\in X$ such that $u_{x}$ and $v_{x}$ have, respectively, the maximum and minimum $x$-coordinate among all points in $X$, and $r$ and $s$ have an analogous property in the $y$-coordinate.
@@ 2 Thomassen’s proof of the WJT
Claim 1. The Intermediate JT holds for PL circuits. Suppose that $f\colon I\to\mathbb{R}^{2}$ is a PL circuit. We first prove that the open set $D=\mathbb{R}^{2}\setminus f[I]$ (Proposition 1.8) is disconnected: we define a continuous map $g\colon D\to\mathbb{R}$ such that $g[D]=\{0,1\}$. By Proposition 1.13, the set $D$ is disconnected because $\{0,1\}$ is a disconnected subset of $\mathbb{R}$.
@@
So it remains to prove that $L(f)$ and $R(f)$ are connected sets. We prove that $L(f)$ is connected (the argument for $R(f)$ is similar). We fix a point $c\in L(f)$ and for every point $d\in L(f)$ we define a connected set $C_{d}$ such that $c,d\in C_{d}\subset L(f)$. Then
@@
We fix one of the two orientations of $C[I]$ and consider the partition $D=\mathbb{R}^{2}\setminus C[I]=L(C)\cup R(C)$ of $D$ defined in Claim 1. Let
@@
Each of these nine open sets $U_{i,j}$ is connected by Propositions 1.15 and 1.16 (the intersection graph spans a path), and they are mutually disjoint. Using Proposition 1.18, we obtain nine PL arcs $h_{i,j}\colon[a,b]\to U_{i,j}$ such that $h_{i,j}$ joins $\alpha_{i,j}=h_{i,j}(a)=g_{i,j}(a)$ ($\in\partial B_{i}$) to $\beta_{i,j}=h_{i,j}(b)=g_{i,j}(b)$ ($\in\partial B_{j}$) in $U_{i,j}$. We consider the first intersections of the straight segments $u_{i}\,\alpha_{i,j}$ and $u_{j}\,\beta_{i,j}$ with $h_{i,j}[I]$ (Proposition 1.9) and obtain nine PL arcs $p_{i,j}$ such that $p_{i,j}$ joins $u_{i}$ to $u_{j}$ and
@@ 3 Filippov’s proof of the WJT
We prove that the parity map $N$ is continuous on $S(f)$. Let $c\in S(f)$. We take a $\delta>0$ small enough such that the ball $B=B(c,\delta)\subset S(f)$ (Proposition 1.8) and that no vertical line that intersects $B$ contains a corner of $f[I]$ distinct from those whose preimages appear in $P(c)$. Let $c^{\prime}\in B$ with $c^{\prime}\neq c$. As we move $c$ to $c^{\prime}$, every simple component of $P(c)$ is preserved and every double component either does not change or disappears or decays in two simple (point) components. It follows that $N(c)=N(c^{\prime})$. So $N$ is continuous at $c$ and on $S(f)$. We obtain the main result of Step 1:
@@
Here is an outline of the proof of this equality: if $S_{i}$ is the set of simple components in $P(c,f_{i})$, $i=1,2,3$, then there is a bijection
$\beta\colon S_{3}\to(S_{1}\times\{0\})\cup(S_{2}\times\{1\})\,.$
We leave the details of $\beta$ to the reader as an exercise.
@@
Step 3. Some two points in the complement of any circuit are separated. Let $f\colon I\to\mathbb{R}^{2}$ be a circuit and $D=\mathbb{R}^{2}\setminus f[I]$. We may assume that $f[I]$ lies to the right of the $y$-axis (Propositions 1.6 and 1.7). Let $l\in f[I]$ and $p\in f[I]$ be, respectively, a leftmost and a rightmost point in $f[I]$ (Propositions 1.6 and 1.19). These points divide $f$ in two arcs $f_{1},f_{2}\colon I\to\mathbb{R}^{2}$ such that $f_{1}^{0}\cap f_{2}^{0}=\emptyset$, $f_{1}[I]\cup f_{2}[I]=f[I]$ and $\mathrm{ep}(f_{1})=\mathrm{ep}(f_{2})=\{l,p\}$ (Proposition 1.21). Clearly, $l_{x}<p_{x}$ (Proposition 1.20). Let $\ell$ be any vertical line with $x$-coordinate in $(l_{x},p_{x})$. Let $a\in f[I]\cap\ell$ be the highest point of this intersection (Proposition 1.19). We may assume that $a\in f_{1}^{0}$. Similarly, let $b\in f_{1}^{0}\cap\ell$ be the lowest point of this intersection (Proposition 1.19). We take a point $c\in\ell$ below $b$ but such that